I need help factoring x^6 - y^6. How would I start?



Found 3 solutions by Earlsdon, khwang, rapaljer:
Answer by Earlsdon(6294)

You could start by recognising that this is the difference of two squares, and recalling that (a^2 - b^2) = (a + b)(a - b) Applying this to the problem at hand:
(x^3)^2 - (y^3)^2 = (x^3 + y^3)(x^3 - y^3)

Answer by khwang(438)

(x^6 - y^6)
= (x^3 + y^3)(x^3 - y^3)
then use (x^3 + y^3) = (x+y) (x^2 - xy+^2)
and (x^3 - y^3) = (x -y) (x^2 + xy+^2)
Normally, I don't answer such kinds of simple questions,
but unfortunately, another person gave you incomplete anwser....
Kenny

Answer by rapaljer(4671)

In defense of Earlsdon, his response was really NOT incomplete, since the question posed was "How would I start?", not "Factor completely!"

See my more detailed explanation of difference and sum of cubes earlier this evening, for those who need help with such simple concepts. And thanks also to you Kenny for handling the harder ones.

R^2 at SCC